OpenStudy (anonymous):

What is the value of the missing exponent? (Type only the numerical value of the missing exponent. Do not skip any spaces.) (2x3)?=4x6

OpenStudy (anonymous):

ok forget that

OpenStudy (anonymous):

i need it fast

OpenStudy (anonymous):

if the question is \[(2\times 3)^?=4\times 6\] you cannot do it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

NO*

OpenStudy (anonymous):

if the six was a nine, the answer would be 2

OpenStudy (anonymous):

(2x^3)^? = 4x^6

OpenStudy (anonymous):

if it is really \[6^x=24\] then the approximate answer is 1.7737

OpenStudy (anonymous):

oh that is an X!! the answer i s 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

YEs i am right for once. thank you

OpenStudy (anonymous):

since \((2x^3)^2=4x^6\)

OpenStudy (anonymous):

now you can put back the happy picture

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hahahahahaha maybe soon

OpenStudy (anonymous):

any more before i go to lunch?

OpenStudy (anonymous):

i have a cooler one now

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

go ahead lets knock them out

OpenStudy (anonymous):

2, 134, 000 = 2.134 x 10^?

OpenStudy (anonymous):

thank you by the way

OpenStudy (anonymous):

\(-\) since you have to move the decimal 6 places left

OpenStudy (anonymous):

damn \(-6\) is what i meant

OpenStudy (anonymous):

thought so

OpenStudy (anonymous):

\[2, 134, 000 = 2.134 \times 10^{-6}\]

OpenStudy (anonymous):

What is the standard form of the number written in scientific notation? 3.14 x 10^6

OpenStudy (anonymous):

6 zeros?

OpenStudy (anonymous):

move the decimal six places to the right

OpenStudy (anonymous):

no, not six zeros, six decimal places

OpenStudy (anonymous):

in between the 2 and 1

OpenStudy (anonymous):

\(3,140,000\)

OpenStudy (anonymous):

the decimal is between the 1 and the 4, if you move it six places to the right you get \[3,140,000.\]

OpenStudy (anonymous):

thank you. two more please

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

by the way, the phone call went well. she felt stupid

OpenStudy (anonymous):

lol she should, that work sheet was a real piece of sheet

OpenStudy (anonymous):

When you write the given number in scientific notation, what is the exponent in the power of 10? 0.000716 = 7.16 x 10^?

OpenStudy (anonymous):

hahahahaha "sheet"

OpenStudy (anonymous):

you have moved the decimal for places to the right, so it is \(-4\)

OpenStudy (anonymous):

yes!!! i dont suck at this, my answers were right